# bibdata.bib

### v1

@@ Abstract
Approximate computing is an effective computing paradigm for improving energy efficiency of error-tolerant applications. Approximate logic synthesis (ALS) is an automatic process to generate approximate circuits with reduced area, delay, and power, while satisfying user-specified error constraints. This paper focuses on ALS under the maximum error constraint. As an essential error metric that provides a worst-case error guarantee, the maximum error is crucial for many applications such as image processing and machine learning. This work proposes an efficient simulation-guided ALS flow that handles this constraint. It utilizes logic simulation to 1) prune local approximate changes (LACs) with large errors that violate the error constraint, and 2) accelerate the SAT-based LAC selection process. Furthermore, to enhance scalability, our ALS flow iteratively selects a set of promising LACs satisfying the error constraint to improve the efficiency. The experimental results show that compared with the state-of-the-art method, our ALS flow accelerates by 30.6×30.6$\times$30.6 ×, and further reduces circuit area and delay by $18.2\%$ and $4.9\%$, respectively. Notably, our flow scales to large EPFL benchmarks with up to 38540 nodes, which cannot be handled by any existing ALS method for maximum error.
Index Terms:
approximate logic synthesis, approximate computing, maximum error, logic simulation
@@ I Introduction
Approximate computing [han2013approximate] is an emerging low-power design paradigm for error-tolerant applications such as signal processing, data mining, and machine learning [mittal2016survey]. It carefully introduces errors to significantly reduce the hardware cost, while the application-level quality is almost unaffected. Approximate logic synthesis (ALS) is an automatic process to generate approximate circuits [scarabottolo2020approximate]. An ALS tool takes an accurate circuit and user-specified error constraints as inputs and outputs an approximate circuit with smaller area, delay, and power, satisfying the constraints.
@@
To evaluate the accuracy of an approximate circuit, two types of error metrics are utilized, the average error and maximum error [scarabottolo2020approximate]. Average error, such as error rate and mean error distance, measures the average deviation between the outputs of the accurate and approximate circuits, while the maximum error measures the maximum deviation between the outputs of the accurate and approximate circuits over all input patterns. Typical maximum errors include maximum error distance (MaxED) and maximum Hamming distance (MaxHD). Maximum error provides a worst-case guarantee of the error, which is crucial for many applications. For instance, in image processing, even if the average error is low, occasional large errors can lead to visible artifacts. Another example is machine learning, where rare extreme deviations may cause mispredictions. Using maximum error constraints ensures that every computation remains within safe bounds, preserving overall quality and reliability. Considering the importance of maximum error, this paper focuses on the ALS under the maximum error constraint.
@@
Many ALS methods for maximum error constraint have been proposed [venkataramani2012salsa, scarabottolo2018circuit, witschen2022muscat, rezaalipour2023parametrizable], most of which simplify the circuit by applying local approximate changes (LACs). A LAC is a local modification of the circuit. For example, a constant LAC [shin2011new] replaces a signal by a constant 0 or 1, and a SASIMI LAC [venkataramani2013substitute] substitutes a signal by another signal. After generating candidate LACs, an ALS flow estimates the maximum error caused by the LACs. Based on the error estimation results, the ALS flow then identifies LACs can be applied to simplify circuit while satisfying the error constraint. There are two categories of maximum error estimation methods. The first category estimates an upper bound of the maximum error [schlachter2017design, scarabottolo2018circuit, scarabottolo2019partition, scarabottolo2021formal]. However, these methods only support the simple constant LAC [shin2011new], while complex LACs such as the SASIMI LAC [venkataramani2013substitute] that can achieve better approximate circuits are not supported. To handle complex LACs, the second category checks whether the maximum error is within a user-specified bound or not [venkatesan2011macaco, chandrasekharan2016approximation, vcevska2017approximating]. For each LAC, these methods convert the maximum error checking problem for the LAC into a SAT problem, and the SAT solving result determines whether the maximum error caused by the LAC is within the bound or not. However, a new challenge is the massive number of complex LACs in a circuit. The solving of their corresponding numerous SAT problems is time-consuming and limits the scalability of the ALS flow. For example, for a circuit with $N$ nodes, there are $O(N^{2})$ SASIMI LACs [venkataramani2013substitute] that replace a signal by another, and solving the corresponding $O(N^{2})$ SAT problems is impractical for large circuits.
@@
To address the above challenges, we propose an efficient logic simulation-guided ALS flow under the maximum error constraint, which can handle complex LACs and scale to large circuits. Logic simulation has shown its effectiveness in accelerating the SAT solving process in many traditional logic synthesis works such as [lee2021simulation], and this motivates us to leverage simulation to accelerate the maximum error checking process in ALS. Our main contributions are as follows:
@@
The experimental results show that compared with the state-of-the-art method, our ALS flow accelerates by 30.6×30.6$\times$30.6 ×, and further reduces the circuit area and delay by $18.2\%$ and $4.9\%$, respectively. Our method can scale to large EPFL benchmarks with up to 38540 nodes, which cannot be handled by any existing ALS method for maximum error constraint. Our work is open-source and available at https://github.com/changmg/SimALS-MaxError.
The remainder of this paper is organized as follows. Section II reviews the related works. Section III introduces the background. Section IV elaborates the proposed simulation-guided ALS flow under the maximum error constraint. The experimental results are presented in Section V, followed by a conclusion in Section VI.
@@ II-A Approximate Logic Synthesis (ALS) Methods
As mentioned before, the average error and the maximum error are two typical error metrics used in ALS. We discuss the ALS methods by the error constraints they support. Some ALS methods can handle both average and maximum error constraints. For instance, Shin and Gupta [shin2011new] propose to replace a signal in the circuit with a constant 0 or 1, supporting the error rate and MaxED constraint. Chandrasekharan et al. [chandrasekharan2016approximation] propose to approximately rewrite AND-inverter graphs (AIGs), which can handle the error rate, MaxED, and MaxHD constraint.
Some ALS methods focus on the average error constraint. Venkataramani et al. [venkataramani2013substitute] propose a LAC called SASIMI, which substitutes a signal with another in the circuit. Their ALS flow considers the error rate or mean error distance constraint. Wu and Qian [wu2016efficient] propose to delete some literals from the Boolean expression of a node in the circuit, and their ALS flow targets the error rate constraint. Hashemi et al. [hashemi2018blasys] utilize Boolean matrix factorization to approximate sub-circuits, considering the mean error distance or mean relative error distance constraint. Zhou et al. [zhou2018dals] developed a delay-driven ALS flow, supporting the error rate constraint, and their work is extended to support the mean error distance and mean Hamming distance constraints in [meng2023hedals]. Meng et al. [meng2020alsrac] simplify the circuit by approximate resubstitution, and their experiments test the error rate, mean error distance, or mean relative error distance constraint. Wang et al. [wang2023dasals] formulates the ALS problem as a neural architecture search problem, and their ALS flow considers the mean squared error constraint.
Other ALS methods deal with the maximum error constraint. Venkataramani et al. [venkataramani2012salsa] propose an ALS method called SALSA. It identifies the approximate don’t cares based on the maximum error limit, and then converts the ALS problem into a traditional logic synthesis problem with don’t cares. However, SALSA requires explicitly representing the don’t cares using the primary inputs (PIs) of a circuit, which is impractical for complex circuits with many PIs. Scarabottolo et al. [scarabottolo2018circuit] propose an ALS method called circuit carving. It identifies the maximum portion of an accurate circuit that can be removed yet still respecting the MaxED constraint. However, this method does not support complex LACs such as the SASIMI LAC [venkataramani2013substitute], hence limiting the hardware cost reduction. Witschen et al. propose an ALS method called MUSCAT [witschen2022muscat]. It formulates the ALS problem under maximum error constraint as a minimal unsatisfiable subset (MUS) problem, whose solution corresponds to an optimal approximate circuit satisfying the error constraint. However, solving the MUS problem is time-consuming, limiting the scalability of MUSCAT. Rezaalipour et al. [rezaalipour2023parametrizable] propose a LAC called XPAT, which uses a parametrizable template to generate an approximate version of a multiple-input multiple-output sub-circuit, but it also has the scalability issue when handling large circuits. Our work differs from the above works on ALS for maximum error constraint in two aspects: 1) supporting complex LACs for better circuit quality, and 2) using effective simulation-guided techniques to enhance scalability.
@@ II-B Error Estimation Methods for ALS
Error estimation of LACs is a critical step in ALS. On the one hand, it avoids violating the error constraint when applying LACs. On the other hand, a more precise error estimation can guide a better selection of LACs so that the final approximate circuit has a smaller hardware cost [su2022vecbee]. The average error estimation is usually based on Monte Carlo simulation [su2022vecbee, meng2022seals, wang2023accals] or analytical methods [echavarria2020probabilistic]. However, these methods cannot be directly applied for maximum error estimation, since the maximum error measures the worst-case deviation, requiring the consideration of all input patterns.
Two categories of methods are proposed to estimate the maximum error caused by LACs. The first category estimates a maximum error upper bound. For example, Schlachter et al. [schlachter2017design] provide a MaxED bound for each node by summing the significance of all reachable outputs of the node. The MaxED bound of a LAC is approximated by the MaxED bounds of the nodes it affects. Scarabottolo et al. [scarabottolo2021formal] propose a partition and propagation technique that derives a tighter maximum error bound for each LAC. However, both [schlachter2017design] and [scarabottolo2021formal] only consider a simple LAC that replaces a signal by a constant 0 or 1 [shin2011new]. Complex LACs such as the SASIMI LAC that can further simplify the circuit are not supported.
The second category of methods directly checks whether the maximum error caused by each LAC is within the error bound. For instance, Venkatesan et al. [venkatesan2011macaco] propose a virtual error circuit (a.k.a. error miter) for checking the maximum error of an approximate circuit. The error miter is then converted into a SAT problem, and the SAT solving result determines whether the maximum error is within the error bound or not. In [chandrasekharan2016approximation] and [vcevska2017approximating], the SAT-based method proposed in [venkatesan2011macaco] is used to check the maximum error caused by each LAC in a circuit. However, given the massive number of LACs, the SAT solver will be called for many times, which is time-consuming and limits the scalability of the ALS flow. Instead of solving SAT problems, the MUSCAT ALS flow [witschen2022muscat] encodes the maximum error checking of all candidate LACs into a single MUS problem. The solution of the MUS problem corresponds to an optimal approximate circuit satisfying the maximum error constraint. However, solving the MUS problem is still time-consuming. Although MUSCAT sets a time limit for the MUS solving to enhance scalability, the time limit may lead to suboptimal approximate circuits. Another approach, the MECALS ALS flow [meng2023mecals], converts the maximum error checking for all candidate LACs into a SAT sweeping problem. Unfortunately, SAT sweeping also has a scalability issue, making MECALS unable to handle large circuits in practice.
In our ALS flow, we use logic simulation to accelerate the check of maximum errors caused by numerous LACs, which belongs to the second category of maximum error estimation methods. Unlike the existing works, our method can handle complex LACs and has better scalability.
@@ III Background
This section introduces the basic concepts of logic circuit terminologies, maximum error metrics, error miter, and LACs.
@@ III-B Maximum Error Metrics
$D_{\text{MaxED}}(\bm{y},\bm{\hat{y}})=\left|{\textit{int}}(\bm{y})-{\textit{int}}(\bm{\hat{y}})\right|,$ (2)
@@ III-C Error Miter for Maximum Error Checking
An error miter is an auxiliary circuit to check whether the maximum error of an approximate circuit exceeds a given bound or not [venkatesan2011macaco, wu2019alfans]. As shown in Fig. 1, it consists of an accurate circuit $G$, an approximate circuit $\hat{G}$, a deviation function unit, and a comparator. The accurate and approximate circuits take the same PIs $\bm{x}$, and their corresponding outputs are $\bm{y}$ and $\bm{\hat{y}}$, respectively. The deviation unit computes the deviation function $D(\bm{y},\bm{\hat{y}})$, such as Eq. (2) or Eq. (3). The comparator checks whether $D(\bm{y},\bm{\hat{y}})$ is larger than the bound $B$. If $D(\bm{y},\bm{\hat{y}})>B$, the output of the comparator, $f$, is $1$; otherwise, $f$ is 00.
To check the maximum error of circuit $\hat{G}$, the error miter is converted into a SAT problem. If the solver returns SAT, then there exists a PI pattern $\bm{x}$ causing $f=1$. In this case, we have $D(\bm{y},\bm{\hat{y}})>B$, and hence ${\textit{MaxError}}(\hat{G})>B$. Otherwise, if the solution is UNSAT, then $f$ is always 00. This means that $D(\bm{y},\bm{\hat{y}})\leq B$ over all PI patterns $\bm{x}$, implying ${\textit{MaxError}}(\hat{G})\leq B$.
@@
III-D Local Approximate Changes (LACs)
A LAC is a local modification of a circuit that can simplify the circuit. A LAC is valid if the approximate circuit after applying the LAC satisfies the error constraint; otherwise, it is invalid.
A single-output LAC replaces a single node $n$ with a new node $n^{\prime}$. For instance, a constant LAC [shin2011new] is a single-output LAC that replaces $n$ by $n^{\prime}=0$ or $n^{\prime}=1$. A SASIMI LAC [venkataramani2013substitute] is also a single-output LAC that replaces $n$ by $n^{\prime}=d$ or $n=\bar{d}$, where $d$ can be any node in the circuit that is not a TFO of $n$. The single-output LACs are the most widely used and can efficiently generate low-cost approximate circuits, as reported in many prior ALS works [shin2011new, venkataramani2013substitute, wu2016efficient, meng2024efficient].
@@ IV-A Overview
As shown in Fig. 2, our flow starts by initializing a current approximate circuit $\hat{G}$ as a copy of the accurate circuit $G$. Then, circuit $\hat{G}$ is iteratively simplified in a main loop, indicated by the blue arrows in Fig. 2. Each iteration consists of three key steps. Step 1 generates a set of candidate LACs $L_{{\textit{cand}}}$. Here, the generated LACs can be any single-output LACs (see definition in Section III-D), such as the constant and SASIMI LACs. Since $L_{\textit{cand}}$ usually contains numerous LACs, step 2 prunes the invalid LACs violating the error constraint according to logic simulation results. The set of remaining LACs after pruning is denoted as $L_{\textit{rem}}$. Then, step 3 selects a set of promising LACs from $L_{\textit{rem}}$ and applies them to simplify the current approximate circuit $\hat{G}$, where a promising LAC refers to a LAC whose application significantly reduces circuit area while satisfying the error constraint. If the approximate circuit $\hat{G}$ is successfully simplified compared to the previous iteration, the main loop continues for the next iteration. Otherwise, no more valid LACs exist, and the main loop terminates. Then, traditional logic synthesis is performed to further simplify the circuit $\hat{G}$ without introducing additional errors, producing the final approximate gate netlist $G_{\textit{final}}$.
@@ IV-B Simulation-Guided LAC Pruning
The LAC pruning step filters out some invalid candidate LACs in $L_{\textit{cand}}$ generated in step 1 of our flow and returns a set of remaining LACs, denoted as $L_{\textit{rem}}$. There are usually many LACs in $L_{\textit{cand}}$. For example, a circuit with $N$ nodes has $O(N^{2})$ SASIMI LACs [venkataramani2013substitute] that replace a node with another. If we use the error miter-based method (see Section III-C) to check the maximum error of each candidate LAC in $L_{\textit{cand}}$, then $O(N^{2})$ SAT problems need to be solved, which is impractical for large circuits. Given that simulation is typically much faster than SAT solving, we propose to use simulation to quickly prune the invalid LACs in $L_{\textit{cand}}$ violating the maximum error constraint. After the pruning, the number of LACs to be checked by SAT solving is significantly reduced, thus improving the efficiency of the ALS flow.

### v2

# Simulation-Guided Approximate Logic Synthesis Under the Maximum Error Constraint

@@ Abstract
Approximate computing is an effective computing paradigm for improving the energy efficiency of error-tolerant applications. Approximate logic synthesis (ALS) is an automatic process to generate approximate circuits with reduced area, delay, and power, while satisfying user-specified error constraints. This paper focuses on ALS under the maximum error constraint. As an essential error metric that provides a worst-case error guarantee, the maximum error is crucial for many applications such as image processing and machine learning. This work proposes an efficient simulation-guided ALS flow that handles this constraint. It utilizes logic simulation to 1) prune local approximate changes (LACs) with large errors that violate the error constraint, and 2) accelerate the SAT-based LAC selection process. Furthermore, to enhance scalability, our ALS flow iteratively selects a set of promising LACs satisfying the error constraint to improve efficiency. The experimental results show that compared with the state-of-the-art method, our ALS flow accelerates by $30.6\times$, and further reduces the circuit area and delay by $18.2\%$ and $4.9\%$, respectively. Notably, our flow scales to large EPFL benchmarks with up to 38540 nodes, which remain challenging for existing ALS methods tackling maximum error constraint.
@@ I Introduction
Approximate computing [10] is an emerging low-power design paradigm for error-tolerant applications such as signal processing, data mining, and machine learning [21]. It carefully introduces errors to significantly reduce the hardware cost, while the application-level quality is almost unaffected. Approximate logic synthesis (ALS) is an automatic process to generate approximate circuits [26]. An ALS tool takes an accurate circuit and user-specified error constraints as inputs and outputs an approximate circuit with smaller area, delay, and power, satisfying the constraints.
@@
To evaluate the accuracy of an approximate circuit, two types of error metrics are utilized, the average error and maximum error [26]. Average error, such as error rate and mean error distance, measures the average deviation between the outputs of the accurate and approximate circuits, while the maximum error measures the maximum deviation between the outputs of the accurate and approximate circuits over all input patterns. Typical maximum errors include maximum error distance (MaxED) and maximum Hamming distance (MaxHD). Maximum error provides a worst-case guarantee of the error, which is crucial for many applications. For instance, in image processing, even if the average error is low, occasional large errors can lead to visible artifacts. Another example is machine learning, where rare extreme deviations may cause mispredictions. Using maximum error constraints ensures that every computation remains within safe bounds, preserving overall quality and reliability. Considering the importance of maximum error, this paper focuses on the ALS under the maximum error constraint.
@@
Many ALS methods for maximum error constraint have been proposed [38, 29, 42, 24], most of which simplify the circuit by applying local approximate changes (LACs). A LAC is a local modification of the circuit. For example, a constant LAC [31] replaces a signal by a constant 0 or 1, and a SASIMI LAC [37] substitutes a signal by another. After generating candidate LACs, an ALS flow estimates the maximum error caused by the LACs. Based on the error estimation results, the ALS flow then identifies LACs that can be applied to simplify the circuit while satisfying the error constraint. There are two categories of maximum error estimation methods. The first category estimates an upper bound of the maximum error [30, 29, 27, 28]. However, these methods only support the simple constant LAC [31], while complex LACs such as the SASIMI LAC [37] that can achieve better approximate circuits are not supported. To handle complex LACs, the second category checks whether the maximum error is within a user-specified bound or not [39, 5, 4]. For each LAC, these methods convert the maximum error checking problem for the LAC into a SAT problem, and the SAT solving result determines whether the maximum error caused by the LAC is within the bound or not. However, a new challenge is the massive number of complex LACs in a circuit. The solving of their corresponding numerous SAT problems is time-consuming and limits the scalability of the ALS flow. For example, for a circuit with $N$ nodes, there are $O(N^{2})$ SASIMI LACs [37] that replace a signal by another, and solving the corresponding $O(N^{2})$ SAT problems is impractical for large circuits.
@@
To address the above challenges, we propose an efficient logic simulation-guided ALS flow under the maximum error constraint, which can handle complex LACs and scale to large circuits. Logic simulation has shown its effectiveness in accelerating the SAT solving process in many traditional logic synthesis works such as [14], and this motivates us to leverage it to accelerate the maximum error checking process in ALS. Our main contributions are as follows:
@@
The experimental results show that compared with the state-of-the-art method, our ALS flow accelerates by $30.6\times$, and further reduces the circuit area and delay by $18.2\%$ and $4.9\%$, respectively. Our method scales to large EPFL benchmarks with up to 38540 nodes, which remain challenging for existing ALS methods tackling the maximum error constraint. Our work is open-source and available at https://github.com/changmg/SimALS-MaxError.
The remainder of this paper is organized as follows. Section II reviews the related works. Section III introduces the background. Section IV elaborates the proposed simulation-guided ALS flow under the maximum error constraint. The experimental results are presented in Section V, followed by conclusions in Section VI.
@@ II-A Approximate Logic Synthesis (ALS) Methods
As mentioned before, the average error and the maximum error are two typical error metrics used in ALS. Some ALS methods can handle both average and maximum error constraints, such as [31] and [5], some focus on the average error constraint, such as [37, 44, 12, 46, 19, 16, 41], and others deal with the maximum error constraint, such as [38, 29, 29, 42, 24]. Our work differs from the above works on ALS for maximum error constraint in two aspects: 1) supporting complex LACs for better circuit quality, and 2) using effective simulation-guided techniques to enhance scalability.
@@ II-B Error Estimation Methods for ALS
Error estimation of LACs is a critical step in ALS. The average error estimation is usually based on Monte Carlo simulation [33, 18, 40] or analytical methods [8]. However, these methods cannot be directly applied for maximum error estimation.
To estimate the maximum error caused by LACs, two categories of methods are proposed. The first category estimates a maximum error upper bound, such as [30, 28]. However, both [30] and [28] only consider a simple LAC that replaces a signal by a constant 0 or 1 [31]. Complex LACs such as the SASIMI LAC [37] that can further simplify the circuit are not supported.
The second category directly checks whether the maximum error caused by each LAC is within the error bound, such as [39, 5, 4, 42, 17]. For example, the MUSCAT ALS flow [42] encodes the maximum error checking of all candidate LACs into a single minimal unsatisfiable subset (MUS) problem. The solution of the MUS problem corresponds to an optimized approximate circuit satisfying the maximum error constraint. However, solving the MUS problem is very time-consuming, and MUSCAT does not support complex LACs like the SASIMI LAC [37]. Although MUSCAT sets a time limit for the MUS solving to enhance scalability, the time limit may lead to suboptimal approximate circuits. Another approach, the MECALS ALS flow [17], converts the maximum error checking for all candidate LACs into a SAT sweeping problem. Unfortunately, SAT sweeping also has a scalability issue, making MECALS unable to handle large circuits in practice. Our method, instead, uses logic simulation to accelerate the check of maximum errors caused by numerous LACs, which belongs to the second category of maximum error estimation methods. Unlike the existing works, our method can handle complex LACs and has better scalability.
@@ III-B Maximum Error Metrics
[…]
@@ III-C Error Miter for Maximum Error Checking
[…]
To check the maximum error of circuit $\hat{G}$, the error miter is converted into a SAT problem. If the solver returns SAT, then there exists a PI pattern $\bm{x}$ causing $f=1$. In this case, we have $D(\bm{y},\bm{\hat{y}})>B$, and hence ${\textit{MaxError}}(\hat{G})>B$. Otherwise, if the solution is UNSAT, then $f$ is always 0. This means that $D(\bm{y},\bm{\hat{y}})\leq B$ over all PI patterns $\bm{x}$, implying ${\textit{MaxError}}(\hat{G})\leq B$.
@@ IV-A Overview
As shown in Fig. 2, our flow starts by initializing a current approximate circuit $\hat{G}$ as a copy of the accurate circuit $G$. Then, circuit $\hat{G}$ is iteratively simplified in a main loop, indicated by the blue arrows in Fig. 2. Each iteration consists of three key steps. Step 1 generates a set of candidate LACs $L_{{\textit{cand}}}$. Here, the generated LACs can be any single-output LACs, i.e., LACs whose affected local circuits have only one output, such as constant LACs and SASIMI LACs. Since $L_{\textit{cand}}$ usually contains numerous LACs, step 2 prunes the invalid LACs violating the error constraint according to logic simulation results. The set of remaining LACs after pruning is denoted as $L_{\textit{rem}}$. Then, step 3 selects a set of promising LACs from $L_{\textit{rem}}$ and applies them to simplify the current approximate circuit $\hat{G}$, where a promising LAC refers to a LAC whose application significantly reduces circuit area while satisfying the error constraint. If the approximate circuit $\hat{G}$ is successfully simplified compared to the previous iteration, the main loop continues for the next iteration. Otherwise, no more valid LACs exist, and the main loop terminates. Then, traditional logic synthesis is performed to further simplify the circuit $\hat{G}$ without introducing additional errors, producing the final approximate gate netlist $G_{\textit{final}}$.
@@ IV-B Simulation-Guided LAC Pruning
The LAC pruning step filters out some invalid candidate LACs in $L_{\textit{cand}}$ generated in step 1 of our flow and returns a set of remaining LACs, denoted as $L_{\textit{rem}}$. There are usually many LACs in $L_{\textit{cand}}$. For example, a circuit with $N$ nodes has $O(N^{2})$ SASIMI LACs [37] that replace a node by another. If we use the error miter-based method (see Section III-C) to check the maximum error of each candidate LAC in $L_{\textit{cand}}$, then $O(N^{2})$ SAT problems need to be solved, which is impractical for large circuits. Given that simulating a small subset of all possible input patterns (e.g., $\leq 2^{13}$ patterns in our experiments) is typically much faster than SAT solving, we propose to use simulation to quickly prune the invalid LACs in $L_{\textit{cand}}$ violating the maximum error constraint. After the pruning, the number of LACs to be checked by SAT solving is significantly reduced, thus improving the efficiency of the ALS flow.